OpenStudy (legendofmath):

what is the function for these coordinates- (1,6) (2,8) (3,10) and (4,12)?

OpenStudy (518nad):

hey how is the x changing and how is the y changing

OpenStudy (legendofmath):

x is being added by 1? and y is being added by 2?

OpenStudy (legendofmath):

PLEASE HELP MEEEE

OpenStudy (color):

y=2x+4

OpenStudy (legendofmath):

that doesnt work work

OpenStudy (legendofmath):

sike

OpenStudy (legendofmath):

im so dumb sorry

OpenStudy (color):

those points lie on a straight line=hint

OpenStudy (phi):

*** x is being added by 1? and y is being added by 2? *** yes, that is good. Next, you have to remember when you see a "constant difference" i.e. when x goes up 1, the y always goes up by 2 that is a *line* (if you plotted the points, you would see this... so if you get stuck, think about plotting the points) Once you have the idea it's a line, write down the formula for the equation of a line y = m x + b Now find the slope (i.e. "m") by finding change in y divided by change in x. Here, the "change in y" is how much it changes i.e. 2, and the change in x is 1 so the slope is 2/1 or just 2 the equation is (replace m with 2) y = 2 x + b to find "b", using any point on the line, for example (1,6) which means when x=1, y=6. Use those numbers: 6 = 2*1 + b 6 = 2 + b add -2 to both sides to get 6-2 = 2-2 + b or 4= b finally, we have y = 2x + 4 once we replace the "b" with its value that is the equation of the points.

OpenStudy (naka354):

Let y = ax + b When x = 1 => 6 = a + b x = 2 => 8 = 2a + b We can get a = 8 - 6 = 2, so y = 2x + b. Substitute x = 1 to y = 2x + b at coordinate (1,6). 6 = 2 + b b = 4 We can conclude that the equation is y = 2x + 4